# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v1

@@ Introduction
$\pi_{k}^{t+1}(a|s)$ $=\mathrm{arg}\min_{\pi}\left\{\left\langle\frac{\partial\tilde{\mathbf{V}}_{\pi_{k}^{t},p}(\rho)}{\partial\theta(s,\cdot)},\pi(\cdot|s)\right\rangle+\frac{1}{\eta^{\prime}}D_{\text{KL}}(\pi(\cdot|s),\pi_{k}^{t}(\cdot|s))\right\}$
$=\mathrm{arg}\min_{\pi}\left\{\frac{1}{1-\gamma}d_{\rho}^{\pi_{k}^{t},p}(s)\left\langle\tilde{\mathbf{Q}}_{\pi_{k}^{t},p}(s,\cdot),\pi(\cdot|s)\right\rangle+\frac{1}{\eta^{\prime}}D_{\text{KL}}(\pi(\cdot|s),\pi_{k}^{t}(\cdot|s))\right\}$
$=\mathrm{arg}\min_{\pi}\left\{\left\langle\tilde{\mathbf{Q}}_{\pi_{k}^{t},p}(s,\cdot),\pi(\cdot|s)\right\rangle+\frac{1}{\eta}D_{\text{KL}}(\pi(\cdot|s),\pi_{k}(\cdot|s))\right\}\,,$ (16)
@@
Denoting $T=\sum_{k=1}^{K}t_{k}$, the oracle version (with perfect gradient info) converges at rate $\mathcal{O}(\log(T)/T)$ in value and constraint violation. PMP-PD Zero (oracle version aiming for zero constraint violation) achieves 00 violation with the same convergence rate. Finally, a version with imperfect gradient information, the so-called sample-based version, was shown to converge at a rate $\tilde{\mathcal{O}}(1/T^{1/3})$, where $T$ denotes the total number of samples, thereby improving on model-free results which have a convergence rate $\mathcal{O}(1/T^{1/4})$.
@@
Having reviewed the performance guarantees of PMD-PD, we now derive the Robust Sample-based PMD-PD algorithm for robust constrained MDPs. The algorithm performs robust training based on Transition Mirror Ascent (TMA) (Wang and Petrik,, 2024) over the robust Lagrangian objective (Bossens,, 2024). Using TMA provides a principled way to compute adversarial dynamics while maintaining the transition dynamics within the bounds of the uncertainty sets. Such uncertainty sets allow for rich parametrisation such as Gaussian mixtures, which can estimate rich probability distributions, and entropy-based parametrisations, which are motivated by the optimal form of KL-divergence based uncertainty sets (see Table 7 for these examples). The algorithm takes place in the sample-based setting, where the state(-action) values are approximated from a limited number of finite trajectories. With a regret bound $\tilde{\mathcal{O}}(\frac{1}{T^{1/3}})$ results are in line with the Sample-based PMD-PD algorithm. The resulting algorithm is given in Algorithm 1. While the first subsections develop theory for the oracle setting, where exact values and policy gradients are given, the subsequent subsection (Section 4.4) derives convergence results for the sample-based setting, where values and policy gradients are estimates. The last subsection then proposes a practical implementation.
Algorithm 1 Robust Sample-based PMD-PD (discrete setting)
Inputs: Discount factor $\gamma\in[0,1)$, error tolerance $\epsilon>0$, error tolerance for LTMA $\epsilon_{0}^{\prime}>0$, failure probability $\delta\in[0,1)$, learning rate $\eta=\frac{1-\gamma}{\alpha}$, dual learning rate $\eta_{\lambda}=1.0$, and penalty coefficients $\alpha=\frac{2\gamma^{2}m\eta_{\lambda}}{(1-\gamma)^{3}}$ and $\alpha_{p}>0$.
Initialise: $\pi_{0}$ as uniform random policy, $\lambda_{0,j}=\max\left\{0,-\eta_{\lambda}\hat{V}_{\pi_{0},p_{0}}(\rho)\right\}$, $p_{0}$ as the nominal transition kernel $\bar{p}$.
for $k\in\left\{0,\dots,K-1\right\}$ do
     for $t\in\left\{0,\dots,t_{k}-1\right\}$ do
         for $(s,a)\in\mathcal{S}\times\mathcal{A}$ do
              Estimate the value: e.g. for softmax parametrisation,
         end for
         $\triangleright$ Policy mirror descent
         if  Using Softmax parametrisation then
              $\pi_{k}^{t+1}(a|s)\leftarrow(\pi_{k}^{t}(a|s))^{1-\frac{\eta\alpha}{1-\gamma}}\exp(-\eta\frac{\hat{\mathbf{Q}}_{\pi_{k}^{t},p_{k}}^{\alpha}(s,a)}{1-\gamma})\,~{}\forall(s,a)\in\mathcal{S}\times\mathcal{A}$.
         else
              $\triangleright$ Use general Bregman divergence
              $\theta_{k}^{t+1}\leftarrow\mathrm{arg}\min_{\theta\in\Theta}\eta\left\langle\nabla_{\theta}\tilde{\mathbf{Q}}_{\pi_{\theta},p_{k}},\theta\right\rangle+\alpha B_{d_{\rho}^{\pi_{\theta},p_{k}}}(\theta,\theta_{k})$$\triangleright$ separate value and divergence estimates
              Define $\pi_{k}^{t+1}:=\pi_{\theta_{k}^{t+1}}$.
         end if
     end for
     Define $\pi_{k+1}:=\pi_{k}^{t_{k}}$
     $\triangleright$ Update transition matrix via Lagrangian TMA (error tolerance fixed to $\epsilon_{0}^{\prime}$ or $\epsilon_{k+1}^{\prime}=\gamma\epsilon_{k}^{\prime}$)
     for $t^{\prime}\in\left\{0,\dots,t_{k}^{\prime}-1\right\}$ do
         $\xi_{k}^{t^{\prime}+1}\leftarrow\mathrm{arg}\max_{\xi\in\mathcal{U}}\eta_{p}\left\langle\nabla_{\xi}\tilde{\mathbf{V}}_{\pi_{k},p_{k}^{t^{\prime}}}(\rho),\xi\right\rangle-\alpha_{p}B_{d_{\rho}^{\pi,p_{k}^{t^{\prime}}}}(\xi,\xi_{k}^{t^{\prime}})$$\triangleright$ Monte Carlo or other techniques
         Define $p_{k}^{t^{\prime}+1}:=p_{\xi_{k}^{t^{\prime}+1}}$
     end for
     Define $p_{k+1}:=p_{k}^{t_{k}^{\prime}}$
     $\triangleright$ Augmented update of Lagrangian multipliers
     Generate $M_{V,k+1}$ samples of length $N_{V,k+1}$ based on $\pi_{k+1}$ and $p_{k+1}$ starting from $s_{0}\sim\rho$ (following Lemma 9).
end for
@@ 4.1 General assumptions
Before listing the general assumptions in the theoretical analysis, we require the definition of rectangular uncertainty sets.
@@
Assumption 1 (Bounded costs and constraint-costs).
Costs and constraint-costs are bounded in $[-1,1]$.
This assumption implies that the Lagrangian at any time introduces a factor $\mathcal{O}(m)$ compared to a traditional value function.
Assumption 2 (Sufficient exploration).
This assumption ensures that when sampling from some $\mu$ instead of the initial state distribution $\rho$, any importance sampling corrections with $\mu$ and $d_{\mu}^{\pi}$ are finite. In particular, it ensures that for any transition kernel $p$ and any policy $\pi$ that the mismatch coefficient $M_{p}(\pi):=\left\lVert\frac{d_{\mu}^{\pi,p}}{\mu}\right\rVert_{\infty}$ is finite. This is also useful for robust MDPs since $M:=\sup_{p\in\mathcal{P},\pi\in\Pi}M_{p}(\pi)$ is also finite.
Assumption 3 (Slater’s condition).
There exists a slack variable $\zeta>0$ such that for any $p\in\mathcal{P}$, there exists a policy $\bar{\pi}$ with $V_{\bar{\pi},p}^{j}(\rho)\leq-\zeta$ for all $j\in[m]$.
Slater’s condition implies strict feasibility of the constraints, zero duality gap, and complementary slackness. However, we do not require knowledge of $\zeta$ in our analysis. The condition is slightly more restrictive than in traditional CMDPs due to the requirement of the condition to hold for all $p\in\mathcal{P}$. However, it is a reasonable assumption since it should typically be assumed for primal-dual algorithms on the transition kernel of interest; the worst-case dynamic $p^{*}$ would typically have a smaller slack anyway.
@@
The uncertainty set is $s$-rectangular or $(s,a)$-rectangular (Def. 1) and is convex, such that for any pair of transition kernel parameters $\xi,\xi^{\prime}$, and any $\alpha\in[0,1]$, if $p_{\xi},p_{\xi^{\prime}}\in\mathcal{P}$ then also $p_{\alpha\xi+(1-\alpha)\xi^{\prime}}\in\mathcal{P}$.
The convex set assumption ensures updates to the transition kernel will remain in the interior so long as it is on a line between two interior points, a common assumption for mirror descent which is also needed for TMA. This is satisifed for most common uncertainty sets, e.g. $\ell_{1}$ and $\ell_{2}$ sets centred around a nominal value. The use of rectangularity assumptions allows us to efficiently use performance difference lemmas for transition kernels as well as make use of results for the optimality of TMA.
Note that when using such parametrised transition kernels, it is convenient to formulate an uncertainty set in the parameter space, which we denote as $\mathcal{U}_{\xi}$. A few examples of parametrised transition kernels (PTKs) and their associated parametrised uncertainty sets can be found in Table 7, including the Entropy PTK (Wang and Petrik,, 2024), which presents a parametrisation related to the optimal parametrisation for KL divergence uncertainty sets of the form $\mathcal{P}_{s,a}=\{p:D_{\text{KL}}(p,\bar{p}(\cdot|s,a))\leq\kappa\}$ (Nilim and Ghaoui,, 2005), and the Gaussian mixture PTK (Wang and Petrik,, 2024), which is able to capture a rich set of distributions for each state-action pair.
@@ 4.2 Global optimality of Lagrangian TMA
As shown below, Lagrangian TMA demonstrably solves the maximisation problem in the robust objective (see Eq. 4) by performing mirror ascent on a suitable PTK. Note that the analysis follows closely that of Wang and Petrik, (2024), where the Lagrangian and the weighted Bregman divergence are explicitly separated in this subsection.
@@
The transition kernel has the gradient
$\nabla_{\xi}\mathbf{V}_{\pi,p}(\rho)=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\pi,p}}[\nabla_{\xi}\log p(s^{\prime}|s,a)(\mathbf{c}(s,a)+\gamma\mathbf{V}_{\pi,p}(s^{\prime}))|\pi,p]\,.$ (18)
@@
The result follows from Theorem 5.7 in Wang and Petrik, (2024) by considering the Lagrangian as a value function (see Eq. 5). ∎
@@
Let $\epsilon^{\prime}>0$, let $\pi$ be a fixed policy, let $\lambda\in\mathbb{R}^{m}$ be a fixed vector of multipliers for each constraint, let $p_{t}$ be the transition kernel at iteration $t$, and let $p_{0}$ be the transition kernel at the start of LTMA. Moreover, let $p^{*}=\mathrm{arg}\min_{p\in\mathcal{P}}\mathbf{V}_{\pi,p}(\rho)$, let $M:=\sup_{p\in\mathcal{P},\pi\in\Pi}M_{p}(\pi)$ be an upper bound on the mismatch coefficient, and let $\eta_{p}>0$ be the learning rate and $\alpha_{p}>0$ be the penalty parameter for LTMA. Then for any starting distribution $\rho\in\Delta(\mathcal{S})$ and any $s\in\mathcal{S}$, the regret of LTMA is given by
$\mathbf{V}_{\pi,p^{*}}(\rho)-\mathbf{V}_{\pi,p_{t}}(\rho)\leq\frac{2F_{\lambda}}{t}\left(\frac{M}{(1-\gamma)^{2}}+\frac{1}{\eta_{p}(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p_{0})\right)\,,$ (19)
where $F_{\lambda}$ is defined according to Eq. 12 for the traditional Lagrangian and on Eq. 11 for the augmented Lagrangian. Moreover, if $\eta_{p}/\alpha_{p}\geq\frac{(1-\gamma)}{2F_{\lambda}}B_{d_{\rho}^{\pi,p^{*}}}(p_{*},p_{0})$ and $t\geq 2F_{\lambda}\frac{M+1}{\epsilon^{\prime}(1-\gamma)^{2}}$, an $\epsilon^{\prime}$-precise transition kernel is found such that
$\mathbf{V}_{\pi,p^{*}}(\rho)-\mathbf{V}_{\pi,p_{t}}(\rho)\leq\epsilon^{\prime}\,.$
@@
Following Theorem 5.5 in Wang and Petrik, (2024), it follows for any value function with cost in $[0,1]$ that
$V_{\pi,p^{*}}(\rho)-V_{\pi,p_{t}}(\rho)\leq\frac{1}{t}\left(\frac{M}{(1-\gamma)^{2}}+\frac{\alpha_{p}}{\eta_{p}(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p_{0})\right)\,.$ (20)
Reformulating the Lagrangian as a value function (see Eq. 5), considering the bounds on the absolute value of the constraint-cost Eq. 10, and observing that the bounds on the (un-)augmented Lagrangian lie in $[-F_{\lambda},F_{\lambda}]$ (see Eq. 11 and  12), the result can be scaled to obtain
$\mathbf{V}_{\pi,p^{*}}(\rho)-\mathbf{V}_{\pi,p_{t}}(\rho)$ $\leq 2F_{\lambda}\left(\frac{M}{(1-\gamma)^{2}}+\frac{\alpha_{p}}{\eta_{p}(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p_{0})\right)\,.$
From the settings of $\eta_{p}/\alpha_{p}\geq\frac{(1-\gamma)}{2F_{\lambda}}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p_{0})$ and $t\geq 2F_{\lambda}\frac{M+1}{\epsilon^{\prime}(1-\gamma)^{2}}$, we obtain
$\mathbf{V}_{\pi,p^{*}}(\rho)-\mathbf{V}_{\pi,p_{t}}(\rho)$ $\leq\frac{2F_{\lambda}}{t}\left(\frac{M+1}{(1-\gamma)^{2}}\right)$

### v2

@@ 2.3 Policy optimisation algorithms
$\pi_{k}^{t+1}(\cdot|s)$ $=\mathrm{arg}\min_{\pi\in\Pi}\left\{\left\langle\frac{\partial\tilde{\mathbf{V}}_{\pi_{k}^{t},p}(\rho)}{\partial\theta(s,\cdot)},\pi(\cdot|s)\right\rangle+\frac{1}{\eta^{\prime}}D_{\text{KL}}(\pi(\cdot|s),\pi_{k}^{t}(\cdot|s))\right\}$
$=\mathrm{arg}\min_{\pi\in\Pi}\left\{\frac{1}{1-\gamma}d_{\rho}^{\pi_{k}^{t},p}(s)\left\langle\tilde{\mathbf{Q}}_{\pi_{k}^{t},p}(s,\cdot),\pi(\cdot|s)\right\rangle+\frac{1}{\eta^{\prime}}D_{\text{KL}}(\pi(\cdot|s),\pi_{k}^{t}(\cdot|s))\right\}$
$=\mathrm{arg}\min_{\pi\in\Pi}\left\{\left\langle\tilde{\mathbf{Q}}_{\pi_{k}^{t},p}(s,\cdot),\pi(\cdot|s)\right\rangle+\frac{1}{\eta}D_{\text{KL}}(\pi(\cdot|s),\pi_{k}^{t}(\cdot|s))\right\}\,$, (16)
@@
Denoting $T=\sum_{k=1}^{K}t_{k}$, the oracle version (with perfect gradient info) converges at rate $\mathcal{O}(\log(T)/T)$ in value and constraint violation. PMP-PD Zero (oracle version aiming for zero constraint violation) achieves 0 violation with the same convergence rate for the value. Finally, a version with imperfect gradient information, the so-called sample-based version, was shown to converge at a rate $\tilde{\mathcal{O}}(1/T^{1/3})$, where $T$ denotes the total number of samples, thereby improving on model-free results which have a convergence rate $\mathcal{O}(1/T^{1/4})$.
@@
Having reviewed the performance guarantees of PMD-PD, we now derive the Robust Sample-based PMD-PD algorithm for robust constrained MDPs. The algorithm performs robust training based on Transition Mirror Ascent (TMA) (Wang et al.,, 2024) over the robust Lagrangian objective (Bossens,, 2024). Using TMA provides a principled way to compute adversarial dynamics while maintaining the transition dynamics within the bounds of the uncertainty sets. Such uncertainty sets allow for rich parametrisation such as Gaussian mixtures, which can estimate rich probability distributions, and entropy-based parametrisations, which are motivated by the optimal form of KL-divergence based uncertainty sets (see Table 7 for these examples). The algorithm takes place in the sample-based setting, where the state(-action) values are approximated from a limited number of finite trajectories. With a regret bound $\tilde{\mathcal{O}}(1/T^{1/3})$, results are in line with the Sample-based PMD-PD algorithm. The resulting algorithm is given in Algorithm 2. While the first subsections develop theory for the oracle setting, where exact values and policy gradients are given, the subsequent subsections derive convergence results for the sample-based setting, where values and policy gradients are estimates. In particular, Section 5.3 proposes and analyses an algorithm for approximate TMA and Section 5.4) proposes and analyses Robust Sample-based PMD-PD. The last subsection then proposes a practical implementation.
@@ 5.1 General assumptions
Assumption 1 (Bounded costs and constraint-costs).
Costs and constraint-costs are bounded in [1,1][-1,1].
The boundedness assumption is ubiquitous in the MDP literature although some authors prefer the range $[0,1]$ or some other constant. For our purposes, it implies that all $1+m$ value functions in the Lagrangian are bounded by $\frac{1}{1-\gamma}$.
Assumption 2 (Bounded range of the dual variables).
The dual variables are upper bounded by $B_{\lambda}>0$.
We assume that the dual variables are bounded by a constant, such that also $F_{\lambda}=\mathcal{O}(1)$, which allows bounding the regret of a policy with respect to the optimal Lagrangian in Eq. 4 to be bounded. The variables can take a wider range compared to other primal-dual algorithms due to the use of the augmented Lagrangian instead of clipping. The parameter is expected to be large for domains with highly active, restrictive constraints. The assumption is not particularly strong since it may be that $B_{\lambda}\gg\frac{2}{(1-\gamma)\zeta}$, the clipping bound usually taken for primal-dual variables (Ding et al.,, 2020). While with the modified Lagrangian update, the iterates obtained in the algorithm are always bounder, Assumption 2 additionally allows bounding the maximising dual variables of each iterate. In the sample complexity analysis of Robust Sample-based PMD-PD, the assumption is only required to limit the average regret induced by a mismatch of the dual variable with the optimal ones for that macro-iteration (see Lemma 28. There, the assumption could be removed when the dual variables are, on average, close to their optima of their respective inner problems.
Assumption 3 (Sufficient exploration).
This assumption is common in MDPs. It ensures that when sampling from some $\mu$ instead of the initial state distribution $\rho$, any importance sampling corrections with $\mu$ and $d_{\mu}^{\pi}$ are finite. In particular, it ensures that for any transition kernel $p$ and any policy $\pi$ that the mismatch coefficient $M_{p}(\pi):=\left\lVert\frac{d_{\mu}^{\pi,p}}{\mu}\right\rVert_{\infty}$ is finite (Mei et al.,, 2020). This is also useful for robust MDPs (Wang et al.,, 2024) since
$M:=\sup_{p\in\mathcal{P},\pi\in\Pi}M_{p}(\pi)$ (18)
is also finite. Similarly, via Lemma 9 of Mei et al., (2020), it ensures $U_{p}=\inf_{s\in\mathcal{S},k\geq 1}\pi_{k}(a_{p}^{*}(s)|s)>0$ when following softmax policy gradient, where $a_{p}^{*}(s)$ is the action for state $s$ under the optimal deterministic policy under the transition dynamics $p$. Again, this also finds its use in robust MDPs (Wang et al.,, 2024) where it follows that
$U:=\inf_{p\in\mathcal{P}}U_{p}>0\,.$ (19)
We use Eq. 19 to upper bound the KL divergence (Lemma 24), by further noting the relation to the global optimum $(\pi^{*},p^{*})$:
$0<U=\inf_{p\in\mathcal{P},s\in\mathcal{S},k\geq 1}\pi_{k}(a_{p}^{*}(s)|s)\leq\inf_{s\in\mathcal{S},k\geq 1}\pi_{k}(a_{p^{*}}^{*}(s)|s)=\inf_{s\in\mathcal{S},k\geq 1}\pi_{k}(a^{*}(s)|s)\,.$ (20)
Assumption 4 (Slater’s condition).
There exists a slack variable $\zeta>0$ and (at least one) policy $\bar{\pi}$ with $V_{\bar{\pi},p}^{j}(\rho)\leq-\zeta$ for all $j\in[m]$ and all $p\in\mathcal{P}$.
The condition is an extension of the traditional Slater’s condition for CMDPs. The assumption of Slater points is not strong: since the constraints are based on expected values, they take on real values. Therefore, if it is possible to obtain the constraint-cost of zero, it is almost surely true that one can also obtain a negative constraint-cost (however small the slack variable). The possible non-negativity of instantaneous constraint-costs does not matter since at least a very small positive budget would be allowed.
In traditional CMDPs, Slater’s condition implies strict feasibility of the constraints, zero duality gap, and complementary slackness for any convex optimisation problem, of which the CMDP is an instance due to its equivalence to a linear programming formulation (Altman,, 1999). The assumption of Slater points is common in CMDPs, and it is inherited from (Liu et al., 2021a, ). It is a ubiquitous but often hidden assumption when using dual and primal-dual algorithms, since it is needed to satisfy strong duality, such that $\min_{\pi\in\Pi}\max_{\lambda\geq 0}\mathbf{V}_{\pi,p}(\rho;\lambda)=\max_{\lambda\geq 0}\min_{\pi\in\Pi}\mathbf{V}_{\pi,p}(\rho;\lambda)$. Many constrained RL algorithms do mention the assumption explicitly (Achiam et al.,, 2017; Liu et al., 2021b, ; Ding et al.,, 2020, 2021). Note that we do not require knowledge of the parameter $\zeta$ while other primal-dual algorithms (Ding et al.,, 2020, 2021) need to know the parameter to clip the Lagrangian. Another approach, which assumes a Slater point but also does not require knowledge of $\zeta$ (Germano et al.,, 2024) bounds the dual variables by showing a strong no-regret property for both the primal and the dual regret minimizers. Approaches without Slater’s condition are possible using primal linear programming (Altman,, 1999; Efroni et al.,, 2020). Unfortunately, such techniques are less practical due to requiring estimates of the occupancy measure.
Assumption 4 is slightly more restrictive than in traditional CMDPs due to the requirement of the condition to hold for all $p\in\mathcal{P}$. The condition of having a subset of the policies being able to strictly satisfy all the constraints for all the transition dynamics reflects the RCMDP objective in Eq. 3 since it is presumed that the worst-case dynamic with respect to the constraints is solvable. For our purposes, the assumption ensures each of the individual CMDPs in the uncertainty set can be individually solved without duality gap similar to Liu et al., 2021b and that the dual variable optimum and iterates (see Lemma 25 and Lemma 27) can be bounded by a constant that can be significantly smaller than the $B_{\lambda}$ that bounds the space of dual variables.
Due to the occupancy being a function of an induced transition kernel which is potentially different for each policy, a zero duality gap may be hard to guarantee for RCMDPs even with Slater points (Wang et al.,, 2022). However, our analysis derives regret bounds with respect to the (primal) global RCMDP optimum. Our sample-based analysis provides a regret upper bound with respect to the (primal) global optimum of the RCMDP $(\pi^{*},p^{*},\lambda^{*})$ using Lemma 8, which is derived based on information from the iteration of the algorithm, namely the error tolerance of LTMA, the suboptimality of the dual variables, and the Lagrangian at $(\pi^{*},p_{k},\lambda_{k-1})$. Such errors are then shown to vanish in an average regret analysis.
@@
Before stating the next assumption, we require the definition of rectangular uncertainty sets.
@@
With the above definition, our assumption on the uncertainty set is as follows.
@@
The uncertainty set is $s$-rectangular or $(s,a)$-rectangular (Def. 1) and is convex, such that for any pair of transition kernel parameters $\xi,\xi^{\prime}\in\mathcal{U}_{\xi}$, and any $\alpha\in[0,1]$, if $p_{\xi},p_{\xi^{\prime}}\in\mathcal{P}$ then also $p_{\alpha\xi+(1-\alpha)\xi^{\prime}}\in\mathcal{P}$.
The convex set assumption ensures updates to the transition kernel will remain in the interior so long as it is on a line between two interior points, a common assumption for mirror descent which is also needed for TMA (Wang et al.,, 2024). The assumption is a standard (often hidden) assumption in gradient descent/ascent type algorithms to ensure the parameter remains in the feasible set.
The rectangular uncertainty set assumption is satisfied for most common uncertainty sets, e.g. $\ell_{1}$ and $\ell_{2}$ sets centred around a nominal value. The use of rectangularity assumptions allows us to efficiently use performance difference lemmas for transition kernels as well as make use of results for the optimality of TMA. The rectangularity assumption is at the basis of the proofs of our sample-based analysis, which uses it to bound the error based on an uncertainty budget for an $\ell_{q}$ uncertainty set. While Wang et al., (2024) mentions the policy optimisation part of their algorithm works with non-rectangular uncertainty sets in principle, the TMA algorithm only applies to rectangular uncertainty sets. Proposals for non-rectangular uncertainty sets do exist but the results are generally not so encouraging. Wiesemann et al., (2013) shows that solving the inner loop of both non-rectangular and non-convex uncertainty sets are both NP hard problems. Recent work demonstrates an actor-critic algorithm with an $\mathcal{O}(\epsilon^{-4})$ bound for non-rectangular uncertainty sets (Li et al.,, 2023). Other alternatives include $r$-rectangular uncertainty sets, which could help reduce the number of distinct factors included to compose the uncertainty set. However, recent work shows that $r$-rectangular and non-rectangular frameworks are only tractable if one assumes costs with no next-state dependencies, an assumption which undermines their usefulness (Grand-Clément et al.,, 2024).
Note that when using such parametrised transition kernels, it is convenient to formulate an uncertainty set in the parameter space, which we denote as $\mathcal{U}_{\xi}$. A few examples of parametrised transition kernels (PTKs) and their associated parametrised uncertainty sets can be found in Table 7, including the Entropy PTK (Wang et al.,, 2024), which presents a parametrisation related to the optimal parametrisation for KL divergence uncertainty sets of the form $\mathcal{P}_{s,a}=\{p:D_{\text{KL}}(p,\bar{p}(\cdot|s,a))\leq\kappa\}$ (Nilim and Ghaoui,, 2005), and the Gaussian mixture PTK (Wang et al.,, 2024), which is able to capture a rich set of distributions for each state-action pair.
@@ 5.2 Global optimality of Lagrangian TMA
As shown below, Lagrangian TMA demonstrably solves the maximisation problem in the robust objective (see Eq. 4) by performing mirror ascent on a suitable PTK. Note that the analysis follows closely that of Wang et al., (2024), where the Lagrangian and the weighted Bregman divergence are explicitly separated in this subsection.
@@
Let $\mathbf{V}_{\pi,p}(\rho;\lambda)$ be the Lagrangian for a policy $\pi\in\Pi$, a transition kernel $p\in\mathcal{P}$ parametrised by $\xi$, and dual variables $\lambda\in\mathbb{R}^{m}$. Then the transition kernel has the gradient
$\nabla_{\xi}\mathbf{V}_{\pi,p}(\rho;\lambda)=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\pi,p}}[\nabla_{\xi}\log p(s^{\prime}|s,a)(\mathbf{c}(s,a,s^{\prime})+\gamma\mathbf{V}_{\pi,p}(s^{\prime};\lambda))|\pi,p]\,.$ (21)
@@
The result follows from Theorem 5.7 in Wang et al., (2024) by considering the Lagrangian as a value function (see Eq. 5). ∎
@@
Let $\epsilon^{\prime}>0$, let $\pi$ be a fixed policy, let $\lambda\in\mathbb{R}^{m}$ be a fixed vector of multipliers for each constraint, let $p^{t}$ be the transition kernel at iteration $t$, and let $p^{0}$ be the transition kernel at the start of LTMA. Moreover, let $p^{*}=\mathrm{arg}\min_{p\in\mathcal{P}}\mathbf{V}_{\pi,p}(\rho;\lambda)$, let $M:=\sup_{p\in\mathcal{P},\pi\in\Pi}M_{p}(\pi)$ be an upper bound on the mismatch coefficient as in Eq. 18, and let $\eta_{p}>0$ be the learning rate and $\alpha_{p}>0$ be the penalty parameter for LTMA. Then for any starting distribution $\rho\in\Delta(\mathcal{S})$ and any $s\in\mathcal{S}$, the regret of LTMA is given by
$\mathbf{V}_{\pi,p^{*}}(\rho;\lambda)-\mathbf{V}_{\pi,p^{t}}(\rho;\lambda)\leq\frac{2F_{\lambda}}{t}\left(\frac{M}{(1-\gamma)^{2}}+\frac{\alpha_{p}}{\eta_{p}(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{0})\right)\,$, (22)
where $F_{\lambda}$ is defined according to Eq. 12 for the traditional Lagrangian and on Eq. 11 for the augmented Lagrangian. Moreover, if $\eta_{p}/\alpha_{p}\geq(1-\gamma)B_{d_{\rho}^{\pi,p^{*}}}(p_{*},p_{0})$ and $t\geq 2F_{\lambda}\frac{M+1}{\epsilon^{\prime}(1-\gamma)^{2}}$, an $\epsilon^{\prime}$-precise transition kernel is found such that
$\mathbf{V}_{\pi,p^{*}}(\rho;\lambda)-\mathbf{V}_{\pi,p^{t}}(\rho;\lambda)\leq\epsilon^{\prime}\,.$
@@
Following Theorem 5.5 in Wang et al., (2024), it follows for any value function with cost in $[0,1]$ that
$V_{\pi,p^{*}}(\rho)-V_{\pi,p^{t}}(\rho)\leq\frac{1}{t}\left(\frac{M}{(1-\gamma)^{2}}+\frac{\alpha_{p}}{\eta_{p}(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{0})\right)\,.$ (23)
Reformulating the Lagrangian as a value function (see Eq. 5), considering the bounds on the absolute value of the constraint-cost Eq. 10, and observing that the bounds on the (un-)augmented Lagrangian lie in $F_{\lambda},F_{\lambda}]$ (see Eq. 11 and  12), the result can be scaled to obtain
$\mathbf{V}_{\pi,p^{*}}(\rho;\lambda)-\mathbf{V}_{\pi,p^{t}}(\rho;\lambda)$ $\leq\frac{2F_{\lambda}}{t}\left(\frac{M}{(1-\gamma)^{2}}+\frac{\alpha_{p}}{\eta_{p}(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p_{0})\right)\,.$
From the settings of $\eta_{p}/\alpha_{p}\geq(1-\gamma)B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p_{0})$ and $t\geq 2F_{\lambda}\frac{M+1}{\epsilon^{\prime}(1-\gamma)^{2}}$, we obtain
$\mathbf{V}_{\pi,p^{*}}(\rho;\lambda)-\mathbf{V}_{\pi,p^{t}}(\rho;\lambda)$ $\leq\frac{2F_{\lambda}}{t}\left(\frac{M+1}{(1-\gamma)^{2}}\right)$